OpenStudy (anonymous):

Lyle has an eleven-cup box of cereal. If one serving is three-fourths cup, how many servings are in the box? Answer 8 and one-fourth eleven and three-fourths fourteen and two-thirds twelve and one-third

OpenStudy (goformit100):

eleven and three-fourths

OpenStudy (anonymous):

11 / (3/4) = 44/3 = 14 + 2/3

OpenStudy (anonymous):

b is the answer